OpenStudy (anonymous):

evaluate 8^2 x 8^-2 not sure what it was asking.... the question was actually like this a^2 x a^-2 then 8=a so yeah i need it evaluated

OpenStudy (anonymous):

can someone walk me through it?

OpenStudy (anonymous):

\[a^2\times a^{-2}=a^{2-2}=a^0=1\]

OpenStudy (anonymous):

or \[a^2\times a^{-2}=\frac{a^2}{a^2}=1\]

OpenStudy (anonymous):

i see thanks :D